OpenStudy (anonymous):

Using complete sentences, explain how to factor 2x2 + 13x + 15. Be sure that the final factorization (or "answer") is a part of your explanation.

jimthompson5910 (jim_thompson5910):

what's the first term?

OpenStudy (anonymous):

2x^2

jimthompson5910 (jim_thompson5910):

what's the coefficient of that term

OpenStudy (anonymous):

2 ?

jimthompson5910 (jim_thompson5910):

what's the last term

OpenStudy (anonymous):

15

jimthompson5910 (jim_thompson5910):

you multiply the first coefficient (2) with the last term (15) to get 2*15 = 30

jimthompson5910 (jim_thompson5910):

now you must find two numbers that a) multiply to 30 AND b) add to 13

OpenStudy (anonymous):

10 * 3 = 30 and 10 + 3 = 13

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that means you have to break 13x into 10x+3x

jimthompson5910 (jim_thompson5910):

2x^2 + 13x + 15 2x^2 + 10x + 3x + 15 now you factor by grouping

jimthompson5910 (jim_thompson5910):

do you know how to do that?

OpenStudy (anonymous):

No thats what I do not get

jimthompson5910 (jim_thompson5910):

ok

jimthompson5910 (jim_thompson5910):

you just pair the terms up like this 2x^2 + 10x + 3x + 15 (2x^2 + 10x) + (3x + 15)

jimthompson5910 (jim_thompson5910):

then you factor out the GCF from each group to get (2x^2 + 10x) + (3x + 15) 2x(x + 5) + 3(x + 5)

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

you have to multiply?

OpenStudy (anonymous):

(2x^2 + 10x) + (3x + 15) 2x(x + 5) + 3(x + 5) 2x + 10 + 3x + 15 ?

jimthompson5910 (jim_thompson5910):

no don't distribute back

jimthompson5910 (jim_thompson5910):

factor further

OpenStudy (anonymous):

what would I factor?

jimthompson5910 (jim_thompson5910):

the two terms are 2x(x + 5) and 3(x + 5)

jimthompson5910 (jim_thompson5910):

what's common between them

OpenStudy (anonymous):

the 5

OpenStudy (anonymous):

and x

jimthompson5910 (jim_thompson5910):

so the x+5

jimthompson5910 (jim_thompson5910):

you would factor that out to get (2x+3)(x+5)

OpenStudy (anonymous):

thats the answer?

jimthompson5910 (jim_thompson5910):

yes 2x^2 + 13x + 15 factors to (2x+3)(x+5)

jimthompson5910 (jim_thompson5910):

expand out (2x+3)(x+5) for yourself and you'll see it simplifies to 2x^2 + 13x + 15

OpenStudy (anonymous):

Can you check an answer for me?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

5x2 − 2x − 7 First multiply 5 and -7. 5 * -7 = -35 Then get a factor of -35 where the sum will be -2. -7 * 5 = -35 and -7 + 5 = 2 Then you will put -7x + 5x where -2x is. 5x^2 − -7x + 5x − 7 Then group the first two terms and you have to pull out x and group the last two terms and factor out 1. 5x2 − -7x + 5x − 7 x( 5x - 7) + 1( 5x - 7) Then you will pull out 5x- 7 because that is the common factor and your answer is (5x - 7) ( x + 1)

jimthompson5910 (jim_thompson5910):

you have two minus signs next to each other in 5x^2 − -7x + 5x − 7

OpenStudy (anonymous):

oh okay thanks other than that?

jimthompson5910 (jim_thompson5910):

other than that, it looks great

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

thanks you so much!

jimthompson5910 (jim_thompson5910):

you're welcome